Operations Report – December 4th

Crew 200 Operations Report 04-Dec-2018

SOL: 3

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart, generator functioning Notes on non-nominal systems: generator was not working this morning (SOC 61% at 8:00), I did not press strong enough the transfer button in the science dome yesterday (did not connect the generator to the power system). Given solar panels were giving enough power, we did not restarted the generator before night (SOC reached 100% during the afternoon).

Generator (hours run): From 5:50 pm last night but stopped at some point during the night. On tonight at 5:00 pm.

Solar— SOC 91 %
Diesel Reading – N/A %
Propane Reading – N/A %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 450 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – not noted gallons, low.
Water (loft) – Static to Loft Pump used – yes
Water Meter: 9

Toilet tank emptied: yes

Deimos rover used: no, still not functional
Hours:
Beginning charge:
Ending charge:
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Opportunity rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Curiosity rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Notes on rovers: Rovers were not used or moved since Sunday, in charge from Sunday afternoon.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: Due to snow and mud conditions, EVA and engineering tour were not allowed. This was approved by Shannon. Thus some of the measurements could not be performed.

Summary of internet: Everything is fine. To avoid using data during the day (indirectly via phone connections), we chosed to switch off the modem when not needed.

Summary of suits and radios: I found out what was wrong with the suits, it is not the suits but the wall electric plugs (the right wall plugs when coming from the indoor). The plugs are not working, thus the chargers were not working. I charged them on another wall plug but it will be needed to have a look at the wall plugs not working. 1 radio seems not to be working.

Summary of Hab operations: we performed first rescue training together this morning. Regular activities went on.

Summary of GreenHab operations: Lights stopped sooner than midnight, permission is asked to fix the clock.

Summary of ScienceDome operations: I looked in detail at the microscopes, I confirmed they are working.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal.

Summary of health and safety issues: Fire extinguishers were transfered for maintenance, Shannon took them.

Questions, concerns and requests to Mission Support:
To Peter: can we access the weather station data? This would be very useful for my energy modelling project to be able to have daily environmental data (especially temperature)
For “water meter” (loft tank), do you mean what is the delta (i.e how much was refilled according to this scale) or the current level? Thanks. [end]